42. You hold a diversified portfolio consisting of an equal investment in each of 15 different common stocks. The portfolio beta is equal to 1.61. You have decided to sell all shares of one of your stocks. The stock you are selling has a beta of 0.45. You plan to use the proceeds to purchase another stock that has a beta equal to 2.70. What will be the beta of the new portfolio?
Answer #1

Increase in beta as a result of net sales=(2.7-0.45)=2.25

Hence increase/stock=(2.25/15)=0.15

Hence beta of new portfolio=(1.61+0.15)

=1.76